OpenStudy (anonymous):

Solve the following system using the substitution method. x-3y=-5 5x+6y=-19

OpenStudy (anonymous):

can you solve for x in the first equation?

OpenStudy (anonymous):

after you do that, plug that expression in for x into the second equation...

OpenStudy (anonymous):

x=3y-5 ?

OpenStudy (anonymous):

good now substitute that expression in for x into the second equation: 5x + 6y = -19 5(3y-5) + 6y = -19 can you solve from here?

OpenStudy (anonymous):

x=-5+3y; 5(-5+3y)+6y=-19; -25+15y+6y=-19; 21y=6; y=6/21...

OpenStudy (anonymous):

That's as far as i got ..

OpenStudy (anonymous):

use that value for y in the first equation to know x

OpenStudy (anonymous):

ok... do you see the post right above^^^ ? that answer is correct for y... all you need to do is use that value for y=6/21 and plug that back into any equation you started with to find the value for x...

OpenStudy (aliahmed):

x=-4.1428 y=0.285714

OpenStudy (anonymous):

Im sooooo lost right now ! This is what my answer options are..

OpenStudy (anonymous):

x=-5+3y now put this in other equation and solve for y!

OpenStudy (anonymous):

once u got the y value put it back in x-3y=-5 and solve for x.

OpenStudy (anonymous):

5x+6y=-19 5(-5-3y)+6y=-19 -25-15y+6y=-19 -9y=6 y=-6/9 y=-2/3 u made the mistake in ur question.. its x+3y=-5* not x-3y=-5

OpenStudy (anonymous):

now to solve for x put y=-2/3 in x+3y=-5 x+3(-2/3)=-5 x-2=-5 x=-5+2 x=-3 so its option 1!!